You are given the following information about an economy:
Gross Investment $40
Govt. purchases of goods & service 30
GNP 200
X – M - 20
Personal Tax 60
Govt. transfer 25
Interest payments from the Govt.
to domestic Pvt. Sector 15
Factor income received from the rest of the would 7
Factor payment made to rest of would 9
Calculate:
a. Consumption b) GDP c) Net factor payment from abroad
b. Pvt. Saving e) Public Saving.
1
Expert's answer
2016-11-04T12:12:08-0400
a. Consumption = C = GNP – (Govt. transfer + Interest payments from the Govt. to domestic Pvt. Sector) = 200 – (25 + 15) = $160
b. GDP = C + I + G + (X - M) = 160 + 40 + 30 – 20 = $210
c. Net factor payment from abroad (NFIA) = Factor income received from the rest of the world – Factor payment made to rest of world = 7-9 = - $2
d. Pvt. Saving = GDP – Personal Tax – Consumption = 210 – 60 – 160 = -$10
e. Public Saving = Personal Tax - Govt. purchases of goods & service = 60-30 = $30